Two random samples are taken, one from among first-year students and the other from among fourth-year students at a public university. Both samples are asked if they favor modifying the student Honor Code. A summary of the sample sizes and number of each group answering yes'' are given below: First-Years (Pop. 1): Fourth-Years (Pop. 2): n 1 =84, n 2 =87, x 1 =34 x 2 =37 First-Years (Pop. 1):n1=84,x1=34Fourth-Years (Pop. 2):n2=87,x2=37 Is there evidence, at an α=0.055 α=0.055 level of significance, to conclude that there is a difference in proportions between first-years and fourth-years? Carry out an appropriate hypothesis test, filling in the information requested (to at least three decimal places, without rounding in intermediate steps). Remember that for the test of equality of two proportions, the SE requires the use of the pooled proportion. A. The value of the standardized test statistic: B. The p-value is D. Your decision for the hypothesis test: A. Do Not Reject H 1 H1 . B. Do Not Reject H 0 H0 . C. Reject H 0 H0 . D. Reject H 1 H1 .
Answer #1

First Years (Group 1)

n1 = 84

p1 = 34/84 = 0.405

Std Dev(s1) = sqrt(p1*(1-p1)/n1) = 0.054

Fourth Years (Group 2)

N2 = 87

P2 = 37/87 = 0.425

Std Dev(s2) = sqrt(p2*(1-p2)/n2) = 0.053

Pooled Variance(sp2) = {(n1-1)s12 + (n2-1)s22}/((n1 + n2 – 2) = 0.003

Hence, Sp = 0.053

Alpha = 0.055

Null and Alternate Hypothesis

H0: µ1- µ2 = 0

Ha: µ1- µ2 <> 0

Test Statistic (t) = (p1-p2)/sp = (0.405-0.425)/0.053 = -0.39

p-value = TDIST(0.39,169,2) = 0.7

Since the p-value is greater than 0.055, we fail to reject the null hypothesis.